OpenStudy (anonymous):

Solve for x: -4|x+5| = -16

OpenStudy (anonymous):

isolate the absolute value to get x+5=4 and x+5=-4 x=-1 and x= -9

OpenStudy (anonymous):

I am sorry i forgot the answers are: x= 21/4, x = -11/4 x = -1 , x = 9 x = -1, x= -9 and no solution

OpenStudy (anonymous):

Thank you masonman